OpenStudy (anonymous):

a 3 ft vertical post casts a 24 inch shadow at the sam time a pine tree casts a 30 ft shadow. How tall is the pine tree?

jimthompson5910 (jim_thompson5910):

Note: 24 inches = 2 feet (since 24/12 = 2) (height of post)/(length of post's shadow) = (height of tree)/(length of tree's shadow) 3/2 = x/30 3*30 = 2*x ... cross multiply and I'll let you finish up

OpenStudy (anonymous):

@jim_thompson5910 thank you

jimthompson5910 (jim_thompson5910):

you're welcome

jimthompson5910 (jim_thompson5910):

tell me what you got so I can check